OpenStudy (anonymous):

if A+B =225, prove that : cot A /1+cot A , Cot b/1+cot B =1/2

OpenStudy (anonymous):

cot A /1+cot A???

OpenStudy (anonymous):

if cotA/(1+cotA) =1/2 cosA/(sinA+cosA) =1/2 tanA=1 similarly tanB=1